Determine the volume of the solid of revolution which is formed when the enclosed are between y=f(x) and y=g(x) is rotated about the y axis.


1
Expert's answer
2020-10-26T20:12:25-0400

By the shell method


"V=\displaystyle\int_{a}^b(2\pi x)\big[f(x)-g(x)\big]dx,"


"f(x)\geq g(x)" on "[a, b]"